# Snail-Radar: A large-scale diverse dataset for the evaluation of 4D-radar-based SLAM systems

### v1

@@ 5 Ground truth
The end subsequence is processed analogously to the start subsequence. However, the end subsequence is reversed to enable processing from the end, as we only know the initial pose at the end. The reverse sequence is created with the requirement that the platform poses remain unchanged in space despite being reversed in time. Specifically, we first find the maximum time of the sequence $t_{max}$, then modify every message’s stamp and every lidar point’s stamp to $t^{\prime}=2t_{max}-t$ from its original time $t$. The accelerometer data remain intact $\mathbf{a}_{m}^{\prime}\leftarrow\mathbf{a}_{m}$, but the gyro data are inverted in sign $\boldsymbol{\omega}_{m}^{\prime}\leftarrow-\boldsymbol{\omega}_{m}$. After processing with LIO in localization mode, we remove the last 8 seconds of the resulting poses for both subsequences. The end subsequence poses’ times are then restored using $t_{max}$. Finally, the start poses and the end poses are combined to form the reference poses for the large scale sequence.
The backward processing also offers a means to verify the reference accuracy by comparing the forward and backward results. For the 20231105/6 sequence, we plot the positions of the lidar relative to the TLS map, $\mathbf{p}_{WL}$ obtained by the forward and backward processing, and their differences $\delta\mathbf{p}_{WL}$ in Fig. 4. The root mean square of $\delta\mathbf{p}_{WL}$ is 9.5 cm, indicating that the reference positions are mostly accurate within a decimeter.
Figure 4: For the 20231105/6 sequence, positions of lidar frames relative to the TLS map obtained through forward and backward processing using the LIO method in localization mode.
@@ 6 Synchronization
Specifically, for the XT32 lidar, we first process the lidar scans with the lidar odometry (LO) method, KISS-ICP (Vizzo et al., 2023), which gives 3D lidar poses at the beginning of each scan. Next, the angular rates in the lidar frame {L} and the INS frame {I} are computed from the INS poses and LO poses by central difference, obtaining $\boldsymbol{\omega}_{WL}^{L}$ and $\boldsymbol{\omega}_{WI}^{I}$. Finally, the time delay of the lidar data is computed by the Correlation Algorithm 1, given the angular rates and the nominal $\mathbf{R}_{IL}$.
For a 4D radar, e.g., the ARS548 or the Oculii Eagle, we first process the radar scans with the GNC method (Yang et al., 2020) to estimate its ego velocity, $\mathbf{v}_{WR}^{R}$. Next, the ego velocity $\mathbf{v}_{WR}^{I}$ is computed from the INS poses with the nominal lever arm $\mathbf{p}_{IR}^{I}$ by central difference. Finally, the time delay of the radar data is estimated using the correlation algorithm, given the ego velocities and the nominal $\mathbf{R}_{IR}$.
For the IMUs, e.g., the ZED2i IMU, which directly provides the angular velocity, $\boldsymbol{\omega}_{WU}^{U}$ in its frame {U}, we first compute the angular rates in the INS frame, $\boldsymbol{\omega}_{WI}^{I}$, by central difference the INS poses, and then run the correlation algorithm given the angular rates and the nominal $\mathbf{R}_{IU}$.
The correlation algorithm 1 finds the time offset $t_{d}$ and relative rotation $\mathbf{R}_{AB}$ between two stamped sequences $\{\mathbf{v}_{i}^{A}\}$ at $\{s_{i},i=1,\cdots,N_{a}\}$ and $\{\mathbf{v}^{B}\}$ at $\{t_{j},j=1,\cdots,N_{b}\}$, e.g., angular rates or ego velocities. Its idea to find the maximum cross correlation is straightforward, but there are a few parameters worth noting. The default smoothing window sizes $M_{a}$ and $M_{b}$ are 1, meaning no smoothing is applied. When the two sequences have very different rates, e.g., 100 vs 10, a smoothing window of size e.g. 5, is necessary for the higher rate sequence. This smoothing causes signal delay which should be accounted for when computing the time offset $t_{d}$. The time tolerance for associating the time-corrected data should typically be smaller than a fraction of the resampling interval, e.g. $\delta t/2$. The truncated least squares method in each iteration first solves for the rotation perturbation $\delta\boldsymbol{\theta}$ and secondly update $\mathbf{R}_{AB}$ by $\mathbf{R}_{AB}\leftarrow\mathrm{Exp}{(\lfloor\delta\boldsymbol{\theta}\rfloor_{\times})}\mathbf{R}_{AB}$. $\delta\boldsymbol{\theta}$ is obtained by solving the linear system stacked from the basic residual equation,
$\lfloor\mathbf{R}_{AB}\mathbf{v}^{B}\rfloor_{\times}\delta\boldsymbol{\theta}\doteq(\mathbf{R}_{AB}\mathbf{v}^{B})\times\delta\boldsymbol{\theta}=\mathbf{v}^{A}-\mathbf{R}_{AB}\mathbf{v}^{B},$ (1)
@@
1:Two sequences of timestamped velocity data $\{\mathbf{v}_{i}^{A}\}$ at $\{s_{i},i=1,\cdots,N_{a}\}$ and $\{\mathbf{v}^{B}\}$ at $\{t_{j},j=1,\cdots,N_{b}\}$; rough guess of the rotation between the two global frames, $\mathbf{R}_{AB}$.
@@
4:Smooth the data by a sliding window of Gaussian weights to remove the high frequency components and improve correlation. The window sizes are $M_{a}$ and $M_{b}$ for the two sequences.
@@
Though the correlation algorithm is robust and globally optimal, it can still fail in certain cases, e.g., when multiple correlation peaks occur. Failed cases are spotted by first checking whether the time offsets along 4 axes (norm, x, y, and z) are consistent and whether the resulting relative rotation is close to the nominal value, and secondly visually checking the time alignment plots. These failures, about 10 out of 300 correlation attempts, were manually corrected case-by-case by tweaking the smoothing window size and/or masking the first peak.
@@
Overall, our sync precision is expected to be better than 10 ms, considering the following facts. 1. The GNSS/INS solution has a rate of 100 Hz, and the correlation algorithm uses quadratic fitting to refine the maximum correlation position. 2. The reported $\sigma$ of Swift-VIO toward the end of a sequence is always less than 1 ms. 3. The exposure time of ZED2i is adjusted to be no more than 5 ms.

### v3

@@ 5.1 TLS-based lidar inertial localization
We also ran the localizer in backward mode to get the backward trajectory, where both the beginning and end subsequences were reversed. The reverse sequence was created with the requirement that the platform poses remained unchanged in space despite being reversed in time. Specifically, we first found the maximum time of the sequence $t_{max}$, then modified every message’s stamp and every lidar point’s stamp to $t^{\prime}=2t_{max}-t$ from its original time $t$. The accelerometer data remained intact $\mathbf{a}_{m}^{\prime}\leftarrow\mathbf{a}_{m}$, but the gyro data were inverted in sign $\boldsymbol{\omega}_{m}^{\prime}\leftarrow-\boldsymbol{\omega}_{m}$. After the processing with LIO in localization mode, we removed the last 8 seconds of the resulting poses for the forward pass of the beginning subsequence and the backward pass of the end subsequence, as these poses were inaccurate. The pose times of the backward pass were then restored using $t_{max}$.
Finally, with the forward and backward trajectories, we took their averages to get the reference trajectory.
The backward pass offers a means to validate the reference trajectory accuracy by comparing the forward and backward results against their averages. For each sequence, we computed the position and rotation difference magnitudes (denoted by $\delta\mathbf{p}_{WL}$ and $\delta\mathbf{R}_{WL}$, respectively) between the forward pass poses and the reference poses. The median and maximum values of these differences are listed in Table 6. Except one trajectory, all sequences have median deviations less than 5 cm and 0.5. Larger deviations typically occur during turns and on bumpy roads.
For the 20231105/4 and 20231105_aft/4 sequences, we plot the position and rotation deviations along the reference trajectory using a jet colormap in Fig. 4. The results show that large deviations are confined to specific locations.
Table 6: The median and maximum differences in translation and rotation between the forward (or backward) trajectory and their averages.
Date/Run Dist(m) $\Delta\mathbf{p}$(cm) $\Delta\mathbf{R}(^{\circ})$
med. max med. max
20230920/1 82.3 0.8 4.5 0.11 3.86
20230920/2 2331.8 1.2 6.3 0.40 2.61
20230921/2 53.5 1.0 4.9 0.11 0.39
20230921/3 2043.6 1.2 9.4 0.17 0.82
20230921/4 716.4 0.9 9.2 0.17 1.83
20230921/5 1976.6 1.0 11.0 0.14 0.75
20231007/2 1983.5 0.9 6.7 0.14 1.78
20231007/4 232.3 0.9 11.4 0.14 1.35
20231019/1 1916.6 1.0 12.1 0.17 2.02
20231019/2 750.9 0.7 9.4 0.15 1.66
20231105/2 2004.3 1.1 32.9 0.16 1.95
20231105/3 2062.8 1.1 15.0 0.14 4.70
20231105/4 897.7 1.4 37.2 0.28 3.99
20231105/5 946.4 1.3 6.9 0.23 1.92
20231105/6 250.1 0.8 7.2 0.13 1.45
20231105_aft/2 166.2 0.7 2.0 0.09 0.39
20231105_aft/4 2000.8 1.2 38.4 0.17 5.29
20231105_aft/5 810.8 0.8 10.5 0.15 1.31
20231109/3 1977.8 1.1 13.9 0.12 1.57
20231109/4 729.4 0.5 4.5 0.05 1.38
20231201/2 1737.3 1.2 10.1 0.04 0.24
20231201/3 1724.3 1.2 7.2 0.04 0.42
20231208/1 275.3 0.6 6.5 0.02 0.19
20231208/4 1718.7 1.2 6.5 0.03 0.20
20231208/5 1731.7 1.3 8.1 0.04 0.29
20231213/1 276.1 1.1 7.0 0.03 0.22
20231213/2 1730.6 1.3 9.3 0.04 0.51
20231213/3 1706.7 1.4 10.3 0.04 0.41
20231213/4 1769.8 1.4 9.0 0.04 0.36
20231213/5 1755.9 1.4 9.0 0.03 0.25
20240113/1 541.4 1.0 10.4 0.03 3.62
20240113/2 1528.1 1.1 7.2 0.03 0.21
20240113/3 1736.9 1.3 6.9 0.03 0.54
20240113/5 1728.1 1.2 6.8 0.03 0.20
20240115/2 1742.2 1.3 8.9 0.03 0.19
20240115/3 1771.9 1.3 8.7 0.03 0.20
20240116/2 648.2 1.1 7.5 0.04 0.24
20240116/4 1706.9 1.4 8.7 0.03 0.20
20240116/5 1744.8 1.2 7.7 0.03 1.74
20240116_eve/3 659.2 1.2 6.6 0.03 0.23
20240116_eve/4 1715.0 1.3 7.1 0.03 0.21
20240116_eve/5 1737.7 1.4 7.2 0.04 0.36
20240123/2 644.6 1.3 7.4 0.04 0.20
20240123/3 1761.9 1.2 8.4 0.03 0.23
(a)     (b)
Refer to caption Refer to caption
    (c)     (d)
Figure 4: The translation (a, c) and rotation (b, d) differences between the forward (or backward) localization and their averages are shown for the e-bike sequences 20231105/4 (a, b) and 20231105_aft/4 (c, d). These two sequences exhibit the largest maximum translation deviations among all sequences in the dataset.
5.2 Cascaded pose graph optimization
To generate full trajectories for sequences extending beyond the TLS map, CPGO (Huai et al., 2018) is applied to fuse multi-source constraints. These constraints include relative poses from the KISS-ICP odometry (Vizzo et al., 2023), absolute poses from TLS-based lidar-inertial localization, absolute positions from the GNSS/INS system, and preintegrated inertial factors from the X36D IMU.
State variables in CPGO consist of the sequential poses and velocities of the XT32 lidar in the TLS map frame, sequential IMU biases, and the gravity direction in the TLS map frame. At first, we solve for the transformation between the UTM50N frame (used for GNSS/INS data) and the TLS map frame by treating it as a variable in CPGO, using a long sequence fully covered by the TLS map (20240921/data5). For other sequences, this transformation is kept constant during CPGO.
CPGO proceeds in three steps. First, it only optimizes the rotation components of the sequential poses by using the relative odometry and absolute TLS poses. Second, it only optimizes the translation components of the sequential poses by using the same constraints. Finally, CPGO refines all state variables by using all available constraints. In this last step, we used the Cauchy loss for relative odometry constraints and GNSS/INS absolute positions to deal with lidar odometry drift in tunnels and GNSS outages. A set of weights for these constraints were manually tuned to improve trajectory agreement across repetitive sessions and used for all sequences. Empirically, we found that cascaded optimization enlarged the convergence basin of the PGO problem. The resulting full trajectories for the XT32 lidar are provided relative to the UTM50N frame. By checking the deviation between two challenging sessions of the August 1 road, we see that the horizontal accuracy of the PGO trajectory is better than 0.15 m, and the vertical accuracy is within 1 meter.
@@ 6 Synchronization
Specifically, for the XT32 lidar, we first processed the lidar scans with the lidar odometry (LO) method, KISS-ICP (Vizzo et al., 2023), which gave 3D lidar poses at the middle of each scan. Next, the angular rates in the lidar frame {L} and the INS frame {I} were computed from the INS poses and LO poses by central difference, obtaining $\boldsymbol{\omega}_{WL}^{L}$ and $\boldsymbol{\omega}_{WI}^{I}$. Finally, the time delay of the lidar data was computed by the Correlation Algorithm 1, given the angular rates and the nominal $\mathbf{R}_{IL}$.
For a 4D radar, e.g., the ARS548 (topic ars548) or the Oculii Eagle (topic radar_pcl2), we first processed the radar scans with the GNC method (Yang et al., 2020) to estimate its ego velocity, $\mathbf{v}_{WR}^{R}$. Next, the ego velocity $\mathbf{v}_{WR}^{I}$ was computed from the INS poses with the nominal lever arm $\mathbf{p}_{IR}^{I}$ by central difference. Finally, the time delay of the radar data was estimated using the correlation algorithm, given the ego velocities and the nominal $\mathbf{R}_{IR}$.
For the IMUs, e.g., the ZED2i IMU, which directly provides the angular velocity, $\boldsymbol{\omega}_{WU}^{U}$ in its frame {U}, we first computed the angular rates in the INS frame, $\boldsymbol{\omega}_{WI}^{I}$, by central differencing the INS poses, and then ran the correlation algorithm given the angular rates and the nominal $\mathbf{R}_{IU}$.
The correlation algorithm 1 finds the time offset $t_{d}$ and relative rotation $\mathbf{R}_{AB}$ between two stamped sequences $\{\mathbf{v}_{i}^{A}\}$ at $\{s_{i},i=1,\cdots,N_{a}\}$ and $\{\mathbf{v}^{B}_{j}\}$ at $\{t_{j},j=1,\cdots,N_{b}\}$, e.g., angular rates or ego velocities. Its idea to find the maximum cross correlation is straightforward, but there are a few parameters worth noting. The default smoothing window sizes $M_{a}$ and $M_{b}$ are 1, meaning no smoothing is applied. When the two sequences have very different rates, e.g., 100 vs 10, a smoothing window of size e.g., 5, is necessary for the higher rate sequence. This smoothing causes signal delay which should be accounted for when computing the time offset $t_{d}$. The time tolerance for associating the time-corrected data should typically be smaller than a fraction of the resampling interval, e.g., $\delta t/2$. The truncated least squares method in each iteration first solves for the rotation perturbation $\delta\boldsymbol{\theta}$ and secondly update $\mathbf{R}_{AB}$ by $\mathbf{R}_{AB}\leftarrow\mathrm{Exp}{(\lfloor\delta\boldsymbol{\theta}\rfloor_{\times})}\mathbf{R}_{AB}$. $\delta\boldsymbol{\theta}$ is obtained by solving the linear system stacked from the basic residual equation,
$\lfloor\mathbf{R}_{AB}\mathbf{v}^{B}\rfloor_{\times}\delta\boldsymbol{\theta}\doteq(\mathbf{R}_{AB}\mathbf{v}^{B})\times\delta\boldsymbol{\theta}\approx\mathbf{v}^{A}-\mathbf{R}_{AB}\mathbf{v}^{B},$ (1)
@@
1:Two sequences of timestamped velocity data $\{\mathbf{v}_{i}^{A}\}$ at $\{s_{i},i=1,\cdots,N_{a}\}$ and $\{\mathbf{v}^{B}_{j}\}$ at $\{t_{j},j=1,\cdots,N_{b}\}$; rough guess of the rotation between the two global frames, $\mathbf{R}_{AB}$.
@@
4:Smooth the data by a sliding window of Gaussian weights to remove the high frequency components and improve correlation. Denote the window sizes by $M_{a}$ and $M_{b}$ for the two sequences.
@@
Though the correlation algorithm is robust and globally optimal, it can still fail in certain cases, e.g., when multiple correlation peaks occur. Failed cases were spotted by first checking whether the time offsets along 4 axes (norm, x, y, and z) are consistent and whether the resulting relative rotation is close to the nominal value, and secondly visually checking the time alignment plots. These failures, about 10 out of 300 correlation attempts, were manually corrected case-by-case by tweaking the smoothing window size and/or masking the first peak.
@@
Overall, our sync precision is expected to be better than 5 ms, considering the following facts. 1. The GNSS/INS solution has a rate of 100 Hz, and the correlation algorithm uses quadratic fitting to refine the maximum correlation position. 2. The reported $\sigma$ of Swift-VIO toward the end of a sequence is always less than 1 ms. 3. The exposure time of ZED2i is adjusted to be no more than 5 ms.
@@ 8 Known issues
9 Comparative studies
Since this dataset is intended for SLAM research, we show its use by evaluating several recent radar-based odometry and place recognition methods, validating its suitability as a benchmarking resource for radar-based SLAM techniques.
9.1 Radar odometry
Recently, numerous radar odometry methods have been proposed. For spinning radars, notable examples include HERO (Burnett et al., 2021) and CFEAR (Adolfsson et al., 2023). For single-chip radars, approaches include radar-only methods such as 4DRadarSLAM (Zhang et al., 2023b) and DeepEgo (Zhu et al., 2023); radar-inertial methods using Doppler data, such as EKF-RIO (Doer and Trommer, 2020) and DRIO (Chen et al., 2023); radar-inertial methods leveraging point matching, such as 4D iRIOM (Zhuang et al., 2023) and (Michalczyk et al., 2022); and learning-based radar-inertial odometry methods like MilliEgo (Lu et al., 2020).
Among them, we selected three recent radar-based odometry methods: 4DRadarSLAM (Zhang et al., 2023b), EKF-RIO (Doer and Trommer, 2020), and 4D iRIOM (Zhuang et al., 2023). These methods were tested on all sequences using data from either the Oculii Eagle or the Continental ARS548 radar. 4D RadarSLAM, specifically designed to process Oculii’s enhanced accumulative point clouds, struggled with the much sparser ARS548 point clouds and the unaccumulated Oculii Eagle data, often failing to find matches. As a result, we evaluated 4DRadarSLAM using the Oculii accumulated point clouds, while the other methods processed the original sparse point clouds.
We measured odometry accuracy using the Absolute Trajectory Error (ATE) metric with the trajectory evaluation tool of Zhang and Scaramuzza (2018). The ATE root mean square errors (RMSEs) are presented in Table 7. Among the methods, 4D iRIOM outperformed the other two across most sequences. However, the recurring large errors underscore that current radar odometry methods are far from robust. While they perform well on select sequences, many sequences exhibit significant drift, indicating substantial room for improvement.
Table 7: Absolute trajectory error (ATE) in meters RMSE (root mean squared error) of radar-based odometry methods, including the 4D RadarSLAM (RS), EKF-RIO (ER), and 4D iRIOM without loop closure (IR), on the Oculii Eagle and Continental ARS548 4D radar data of the proposed dataset. The rows are grouped according to the routes as in Table 2. Note that the ARS548 data are unavailable for sequences before 20231105. The 4D RadarSLAM needs to run with the enhanced accumulative Oculii radar data as it requires dense points for scan matching. The other two methods, EKF-RIO and 4D iRIOM processes the single frame data from Eagle and ARS548 without accumulation for better results. A ’-’ indicates that ARS548 data is not available.
Sequence Eagle ARS548
RS ER IR ER IR
230920/1 1.6 6.0 0.7 - -
230921/2 1.4 6.4 0.9 - -
231007/4 9.5 12.1 4.7 - -
231105/6 6.1 1.6 1.6 8.9 0.7
231105_aft/2 1.8 1.5 3.2 0.8 0.6
230920/2 156.9 188.3 58.9 - -
230921/3 200.3 202.6 134.2 - -
230921/5 174.6 194.7 58.7 - -
231007/2 200.5 194.2 68.4 - -
231019/1 64.7 197.4 58.2 - -
231105/2 37.8 143.9 31.7 137.0 25.8
231105/3 118.3 134.5 62.0 102.8 33.5
231105_aft/4 167.8 125.5 72.4 69.0 9.1
231109/3 190.2 128.3 209.2 61.0 11.6
230921/4 27.5 58.3 12.6 - -
231019/2 35.4 57.6 59.1 - -
231105/4 18.2 25.8 2.4 37.5 17.9
231105/5 26.6 27.9 52.6 39.4 15.1
231105_aft/5 32.1 16.8 51.2 20.9 10.6
231109/4 37.3 31.2 45.7 22.0 16.3
231208/4 75.8 60.7 8.5 43.4 5.7
231213/4 4.9 80.5 30.8 47.4 2.0
231213/5 11.6 64.7 22.7 48.2 2.8
240115/3 25.9 46.9 3.1 52.5 5.0
240116/5 42.4 75.7 32.7 56.4 4.4
240116_eve/5 18.5 57.9 23.3 51.7 3.2
240123/3 58.4 53.6 22.3 53.3 4.2
231201/2 166.5 117.4 41.4 85.0 1.2
231201/3 92.9 53.9 44.5 94.7 9.3
231208/5 129.7 116.5 54.5 28.8 1.0
231213/2 134.5 187.9 2.1 180.1 7.4
231213/3 129.6 210.8 100.5 119.2 23.2
240113/2 184.3 182.1 31.1 175.2 77.1
240113/3 101.8 105.2 77.7 63.8 1.3
240116_eve/4 145.9 186.5 96.0 179.5 11.9
240113/5 171.2 138.3 116.8 58.6 1.5
240115/2 159.7 208.8 79.7 114.7 1.0
240116/4 169.4 170.5 28.0 216.7 1.3
231208/1 2.1 8.1 4.7 5.1 0.6
231213/1 1.7 7.5 1.7 5.9 0.8
240113/1 32.6 15.8 3.2 12.4 0.5
240116/2 69.6 76.6 68.4 78.5 0.2
240116_eve/3 76.5 76.6 63.7 76.9 0.2
240123/2 8.6 73.6 31.3 73.1 0.3
9.2 Radar place recognition
Likewise, numerous radar-based place recognition methods have been developed over the years. For spinning radars, notable approaches include k-Radar++ (De Martini et al., 2020), Kidnapped Radar (Săftescu et al., 2020), Open-RadVLAD (Gadd and Newman, 2024), ReFeree (Kim et al., 2024), and RaPlace (Jang et al., 2023). These methods primarily work with range-azimuth data.
In contrast, 4D single-chip radars provide additional Doppler information and vertical data, offering the potential for improved place recognition. Several learning-based methods utilizing single-chip radars have been proposed, including AutoPlace (Cai et al., 2022) designed for 3D radars, voxel-based TransLoc4D (Peng et al., 2024), and SPR for single sparse scans (Herraez et al., 2024), and 4D RadarPR that adaptively fuses point features and multiscale context features (Chen et al., 2025).
In this work, we evaluate three radar-based place recognition methods using either ARS548 or Eagle radar data from our dataset. These methods include the classic point-based PointNetVLAD (Uy and Lee, 2018), the projection-based AutoPlace (Cai et al., 2022), and TransLoc4D (Peng et al., 2024), all of which were retrained from scratch using each type of radar data in our dataset.
To begin, the dataset is divided into three subsets: training, validation, and test. The training set consists of seven sequences from four routes: Starlake, Software School, Info Faculty, and August 1 Road, captured under various environmental conditions. The validation set, used to assess the model’s effectiveness during training, includes four sequences from both the Info and Arts Faculty and the Info Faculty routes. The remaining sequences, which are not part of the training or validation sets, form the test set. These test set includes some sequences from routes unseen in the training data. In both the validation and test sets, there are multiple sequences for a route. For these, the first sequence serves as the database, while the others act as queries. During training, thresholds for positive and negative samples were set to 9 m and 18 m, respectively, with a 9 m threshold applied for positive samples during testing. Additionally, given the radar’s approximate 120 field of view, the heading angle difference threshold was set to 75 for training and 30 for testing to keep positive candidates.
For data preparation, moving objects were filtered using the GNC method as described in Zhuang et al. (2023). Frames were then selected at intervals of five for place recognition. For the Eagle radar, enhanced accumulative point cloud frames were used. To address the sparsity of the ARS548 point clouds, a window of seven frames centered on the selected frames was used to form an aggregated frame, with compensation for relative motion derived from the reference trajectory.
For all three methods, we primarily use the default settings from the public codebases unless otherwise specified. For PointNetVlad, the input consists of 4D tuples of point coordinates and RCS, and the output is a 256D descriptor. The model was trained for 40 epochs using the lazy quadruplet loss with a learning rate of $1\cdot 10^{-5}$ and a batch size of 2. Each batch includes 1 query frame, 2 positive frames, 18 negative frames, and 1 hard negative frame. For AutoPlace, the model was trained for 200 epochs with an initial learning rate of $1\cdot 10^{-4}$ and a batch size of 16. For TransLoc4D, the input consists of 5D tuples ($\mathbf{p}$, $s$, $I$), where $\mathbf{p}$ is radar point coordinates, $s$ is the relative azimuth angle, and $I$ is the RCS, and the output is also 256D. The model was trained for 500 epochs with an initial learning rate of $1\cdot 10^{-3}$ and a batch size of 768.
The recall rates for the three methods are listed in Table 8. Despite its age, PointNetVlad, originally developed for lidar, often produced the best place recognition results. The relatively low recall in the top five candidates using the Eagle radar may be attributed to motion distortion in the accumulative radar point clouds. The suboptimal performance of the recent TransLoc4D method may be attributed to its reliance on velocity directions for generating a place descriptor, which vary across different traversals (Herraez et al., 2024).
Table 8: Recall of three recent place recognition methods on the proposed dataset, using the Continental ARS548 and Oculii Eagle enhanced 4D radar data. The three methods include the classic point-based PointNetVlad, the projection-based AutoPlace, and the voxel-based TransLoc4D. r@K represents the recall when considering the top K retrieved samples. K can also be set as 1% of the total number of descriptors in the database. The route abbreviations follow Table 2.
Routes Methods Continental ARS548 Oculii Eagle enhanced
r@1 r@5 r@10 r@1% r@1 r@5 r@10 r@1%
bc PointNetVlad 81.0 93.6 98.9 100.0 72.2 82.7 87.6 76.5
AutoPlace 76.8 90.2 93.9 88.6 71.3 83.5 89.6 75.4
TransLoc4D 35.9 74.6 84.7 91.5 49.4 69.5 76.1 84.1
sl PointNetVlad 59.1 79.0 85.7 89.7 64.8 76.5 81.0 86.6
AutoPlace 43.6 60.8 68.0 72.2 47.6 58.6 64.9 80.2
TransLoc4D 21.9 45.8 56.9 70.1 49.2 68.5 75.4 89.5
ss PointNetVlad 98.3 99.8 100.0 100.0 57.9 70.1 76.3 80.2
AutoPlace 97.5 99.7 100.0 100.0 41.2 52.4 59.7 65.7
TransLoc4D 88.8 96.9 97.7 99.0 46.1 65.0 73.6 83.4
if PointNetVlad 97.8 99.4 99.6 99.7 95.9 97.7 98.0 98.2
AutoPlace 95.4 98.3 99.0 99.4 88.7 94.9 96.4 96.6
TransLoc4D 90.7 98.6 99.3 99.6 86.5 96.7 98.1 98.8
iaf PointNetVlad 96.9 99.0 99.3 99.6 92.4 96.6 97.7 98.8
AutoPlace 96.2 98.6 99.2 99.3 85.3 92.4 95.1 97.7
TransLoc4D 91.0 98.1 99.0 99.6 80.5 94.5 97.2 98.8
iaef PointNetVlad 97.3 99.1 99.4 99.7 87.9 93.5 95.4 97.3
AutoPlace 95.6 97.6 98.2 99.2 81.8 89.1 92.7 96.7
TransLoc4D 92.6 98.0 98.6 99.5 73.0 87.3 90.2 94.1
st PointNetVlad 100.0 100.0 100.0 100.0 98.5 99.6 100.0 100.0
AutoPlace 99.7 99.8 99.9 100.0 93.3 96.4 98.0 96.1
TransLoc4D 99.4 100.0 100.0 100.0 91.7 96.3 97.1 98.7
81r PointNetVlad 93.1 97.0 97.8 98.8 87.7 94.7 96.5 97.6
AutoPlace 94.3 98.1 99.0 99.8 85.2 92.3 94.4 96.8
TransLoc4D 89.8 97.5 98.7 99.9 73.4 91.0 95.3 98.7